OpenStudy (anonymous):

How do you derive this velocity function into acceleration? y=-50cos(x)sin(25x)-2sin(x)cos(25x)

OpenStudy (anonymous):

preferably using the product rule

terenzreignz (terenzreignz):

It doesn't look like you have much of a choice but to use the product rule :) But can you also use chain rule?

OpenStudy (anonymous):

yeah we can use whatever rule we like i guess :) i just dont even know how to do it..

terenzreignz (terenzreignz):

Chain rule, though. It's kind of the key here. Can you use it?

OpenStudy (anonymous):

yeah i can, but i dont know where to use it exactly?

terenzreignz (terenzreignz):

Okay... first, we clarify a few things...if y is the velocity function, then the acceleration is the derivative of the velocity function...\[\large a=\frac{d}{dx}\left( -50\cos(x)\sin(25x)-2\sin(x)\cos(25x) \right)\]

OpenStudy (anonymous):

yep, i get that :)

OpenStudy (anonymous):

what do you mean by sin(25x)????

terenzreignz (terenzreignz):

Well, you know that the differentiation operator distributes over addition (subtraction) and that constants can just be moved out (in other words, it's a linear operator) So...

OpenStudy (anonymous):

uh, -50cos(x)*sin(25x)

terenzreignz (terenzreignz):

\[\large a= -50\frac{d}{dx}(\cos(x)\sin(25x))-2\frac{d}{dx}(\sin(x)\cos(25x)) \]

OpenStudy (anonymous):

I have a question, the derivative of sinx is cosx, so is the derivative for sin(25x) = cos(25x)

OpenStudy (anonymous):

is 25cos25x

OpenStudy (anonymous):

whatever is in front of the x value comes out the front also. Im still a little condfused?

terenzreignz (terenzreignz):

Let's focus on this part \[\large \frac{d}{dx}(\cos(x)\sin(25x))\]Product rule... can you state it?

OpenStudy (anonymous):

y=uv'+u'v

OpenStudy (anonymous):

there is nothing called as such, cos(25x) means the angle is only multiplied by 25, however 25cos(x) means that 25 is multiplied by the cos value of x

OpenStudy (anonymous):

what about the constant, -50?

OpenStudy (anonymous):

woah, what? we were taught that whatever value that is in front of x, comes out the front and multiplies with any constants?

terenzreignz (terenzreignz):

The constant was taken out of the differentiation operator. You know that \[\large \frac{d}{dx}cf(x)=c\frac{d}{dx}f(x)\]

OpenStudy (anonymous):

no, whats that?

terenzreignz (terenzreignz):

Constants may move in and out of derivatives :D

OpenStudy (anonymous):

aha okay :)

terenzreignz (terenzreignz):

That is to say, the derivative of a constant times the function is the constant times the derivative of the function.

OpenStudy (anonymous):

True.

OpenStudy (anonymous):

ah okay, so the -50 stays out the front

terenzreignz (terenzreignz):

We can prove that... using the product rule... \[\large \frac{d}{dx}cf(x)=c\frac{d}{dx}f(x)+f(x)\frac{d}{dx}c\]Derivative of a constant is zero... \[\large \frac{d}{dx}cf(x)=c\frac{d}{dx}f(x)+f(x)\cdot0=c\frac{d}{dx}f(x)\]done. Shall we continue?

OpenStudy (anonymous):

sure :)

terenzreignz (terenzreignz):

Back to this... \[\large \frac{d}{dx}(\cos(x)\sin(25x))\] You can let u = cos x and v = sin(25x) Now apply the product rule...

OpenStudy (anonymous):

oh, so you have to do the product rule for each side?

terenzreignz (terenzreignz):

What do you mean by "each side" The other side (left) only has a...

terenzreignz (terenzreignz):

Or did you mean for each term? In that case, yes.

OpenStudy (anonymous):

yeah term. I have a question. -50cos(x) derives to be -50(-sinx), but do the negatives cancel?

terenzreignz (terenzreignz):

Of course they do :) The things you learned in algebra still apply... very much :D

OpenStudy (anonymous):

alright, so u'=50sin(x) and v'=-2cos(x)*25-sin(25x), so 50cos(x)sin(25x)

OpenStudy (anonymous):

y=-1250*(cos x)*cos (25*x) + 50*(sin (25*x))*sin x - (-50*(sin x)*sin (25*x) + 2*(cos (25*x))*cos x) that is the derivative from graphmatica

terenzreignz (terenzreignz):

Nope... the derivative of a product is NOT the product of their derivatives... (uv)' = uv' + u'v You stated this... remember? :)

OpenStudy (anonymous):

so what did i do wrong?

OpenStudy (anonymous):

oh wait, i was doing the other terms!

terenzreignz (terenzreignz):

You made it so that (uv)' = u'v'

terenzreignz (terenzreignz):

too. So careful

OpenStudy (anonymous):

u'=50sin(x) and v'=25cos(x)

terenzreignz (terenzreignz):

Okay. Now product rule. And be careful ;)

terenzreignz (terenzreignz):

Wait... v' Redo it.

OpenStudy (anonymous):

y'=-1250cos(x)cos(25x)+50sin(x)sin(25x)

OpenStudy (anonymous):

yeah i fixed v' up, my mistake!

terenzreignz (terenzreignz):

Wait....patience... let's start over... or at least... why is it 1250?

OpenStudy (anonymous):

constants multiplied. -50cos x 25cos.. -50x25=-1250

OpenStudy (anonymous):

that side is right according to graphmatica :)

terenzreignz (terenzreignz):

But... something's wrong... we already brought out the -50. Why did you include it in your u?

OpenStudy (anonymous):

cause i hadnt bought it in anywhere. I included it with my u because i let u=-50cos(x)

terenzreignz (terenzreignz):

Okay, I've seen my error. Sorry, my bad. Please continue.

OpenStudy (anonymous):

when you do the product rule for the second term, what do you do with them all at the end? and thats alright!

terenzreignz (terenzreignz):

You add them.

terenzreignz (terenzreignz):

Keeping in mind that there was a minus sign in between them, of course.

terenzreignz (terenzreignz):

Remember...\[\large a= -50\frac{d}{dx}(\cos(x)\sin(25x))-2\frac{d}{dx}(\sin(x)\cos(25x))\]And you've already done the first product rule bit...

OpenStudy (anonymous):

so when i do the second product rule bit, i just add them all other than the minus?

terenzreignz (terenzreignz):

Well, distribute the minus sign. Differentiate the second term, first. I can't put in words what to do, I'd rather demonstrate, to avoid confusion.

OpenStudy (anonymous):

hold on, would you be able to check my second product rule part?

terenzreignz (terenzreignz):

Sure

OpenStudy (anonymous):

alright. i let u=-2sin(x), so u'=-2cos(x) i let v = cos(25x), so v'=-25sin(25x). When i used the product rule, i got: 50sin(x)sin(25x) - 2xos(x)cos(25x)

terenzreignz (terenzreignz):

Nicely done :) Now the only thing left to do is add them up :D

OpenStudy (anonymous):

i dont think im going to get the same as the graphing program though..

terenzreignz (terenzreignz):

I know it's daunting to check it, but actually, when you add them up, you DO get the same answer :)

OpenStudy (anonymous):

so it would be: -1250cos(x)cos(25x)+50sin(x)sin(25x)-50sin(x)sin(25x)+2cos(x)cos(25x)

terenzreignz (terenzreignz):

Copying your post... y=-1250*(cos x)*cos (25*x) + 50*(sin (25*x))*sin x - (-50*(sin x)*sin (25*x) + 2*(cos (25*x))*cos x)

OpenStudy (anonymous):

for the second half, it says that it is meant to be -50(-sin...) but, considering that in the product rule it was -2sin(x)-25sin(25x), wouldnt the negatives cancel?

terenzreignz (terenzreignz):

And uhh, you got a sign wrong...

OpenStudy (anonymous):

and the minus sign would just make it -50

terenzreignz (terenzreignz):

In this post... so it would be: -1250cos(x)cos(25x)+50sin(x)sin(25x)-50sin(x)sin(25x)+2cos(x)cos(25x) You got a sign wrong right.............................................^ here

OpenStudy (anonymous):

im confused about the sign too

terenzreignz (terenzreignz):

Because you also posted... alright. i let u=-2sin(x), so u'=-2cos(x) i let v = cos(25x), so v'=-25sin(25x). When i used the product rule, i got: 50sin(x)sin(25x) - 2xos(x)cos(25x) See this sign ......^ :)

OpenStudy (anonymous):

yeah makes sense! why is it + in the answer? and why is it -50(-sin..)

terenzreignz (terenzreignz):

Oh... I see. You distributed the minus sign. But that wasn't necessary, as you already considered the minus sign when you let u = -2sin(x) Signs are VERY confusing, so be EXTRA careful, ok?

OpenStudy (anonymous):

so what have i done wrong? im so confused now :(

OpenStudy (anonymous):

also, can you have y''=? cause its derived twice

terenzreignz (terenzreignz):

What you did was you distributed the minus sign like this -(50sin(x)sin(25x)-2cos(x)cos(25x))When you shouldn't have done so, because, there is no longer a minus sign there, as you have already considered that minus sign when you let u = -2sin(x)

OpenStudy (anonymous):

and i cant have acceleration, cause there is no time!

terenzreignz (terenzreignz):

x is time.

terenzreignz (terenzreignz):

At least, so I assumed.

OpenStudy (anonymous):

a(x) then? and how do i change it so its right?

terenzreignz (terenzreignz):

Again, just add them up, without frills.

terenzreignz (terenzreignz):

Because this has been so long, I'll repost your posts.. y'=-1250cos(x)cos(25x)+50sin(x)sin(25x) alright. i let u=-2sin(x), so u'=-2cos(x) i let v = cos(25x), so v'=-25sin(25x). When i used the product rule, i got: 50sin(x)sin(25x) - 2xos(x)cos(25x)

OpenStudy (anonymous):

so the second half of the product rule would be -50(-sinx)sin(25x)-2cos(x)cos(25)

terenzreignz (terenzreignz):

Yeah. And add them up.

OpenStudy (anonymous):

but why is the sin still negative if it cancels?

terenzreignz (terenzreignz):

It isn't.

OpenStudy (anonymous):

-2sin(x)-25sin(x)-2cos(x)cos(25x) would be: 50sin(x)sin(25x)-2cos(x)cos(25x) cause the -2 and 025 make positive 50. But when you combine it all, the minus goes before the 50, but why does the sin become negative again?

terenzreignz (terenzreignz):

The first time you used the product rule on the second term, you got it right, but then, you incorrectly took its negative when you shouldn't have done so... Posting it again... alright. i let u=-2sin(x), so u'=-2cos(x) i let v = cos(25x), so v'=-25sin(25x). When i used the product rule, i got: 50sin(x)sin(25x) - 2xos(x)cos(25x)

OpenStudy (anonymous):

y=-1250*(cos x)*cos (25*x) + 50*(sin (25*x))*sin x - (-50*(sin x)*sin (25*x) + 2*(cos (25*x))*cos x) ....................................................................................^

terenzreignz (terenzreignz):

sin became negative because the derivative of cos x is -sin x Also, it's 50sin(x)sin(25x) by the way

OpenStudy (anonymous):

convert the cos*cos and sin*"sin products to sums your life'd get much easier

OpenStudy (anonymous):

-2sin(x)*-25sin(25x) = 50sin(x)sin(25x) am i right?

terenzreignz (terenzreignz):

That's right.

OpenStudy (anonymous):

\[\cos(x)\sin(25x)={1\over2}[\sin(x+25x)+sin(x-25x)]\\ \sin(x)\cos(25x)={1\over2}[\sin(x+25x)-\sin(x-25x)] \]

OpenStudy (anonymous):

all you are left with is buch of sines adding

OpenStudy (anonymous):

si why is it that when you combine and add them all together, it goes - 50 (-sin..) where did the (-sin..) come from if the negatives cancelled to make a positive!

terenzreignz (terenzreignz):

Where the -sin came from? Because the derivative of cos x is -sin x